# Privacy Risks and Preservation Methods in Explainable Artificial Intelligence: A Scoping Review

### v2

@@ 5.3 Anonymization
Anonymization refers to the process of transformation of data (Majeed & Lee, 2021) to obscure the distinctive features of individuals, thus safeguarding their privacy. The process is associated with the removal or modification of direct and quasi-identifiers (Majeed & Lee, 2021), that can uniquely identify individuals. Various methods of anonymization are used in practice, such as k-anonymity, l-diversity, and t-closeness (Yin et al., 2022). In XAI, different anonymization techniques have been demonstrated in example-based, feature-based, and interpretable methods. Techniques such as disentangled representation learning and lossy compression have been applied on sensitive visual data, such as medical images, to generate privatized explainable-by-design representations.
@@
A dataset is considered k-anonymous if every record is indistinguishable from k-1 other records (Sweeney, 2002b), thus providing a measure of the risk of re-identification of records. K-anonymity can be achieved using methods such as suppression and generalization (Sweeney, 2002a), which obscure the data to remove identifiable features. Though traditionally this technique is applied to target datasets for protection, Goethals et al. (2023) proposed its usage on native counterfactuals, that are actual datapoints from the training dataset, for protection against model inversion attack through explanations. This strategy of generating k-anonymous counterfactuals was shown to result in lower information loss and higher validity, outperforming counterfactuals derived from k-anonymized datasets. Further analysis by Berning et al. (2024) determined that the effectiveness of k-anonymous counterfactuals is confined to dense areas of the dataset. Its offered privacy protection was also found disproportionate to the value of k. Vo et al. (2023) highlighted another limitation, namely, the computational overhead of generating these counterfactuals requiring querying the explainer for a large number of counterfactuals. The authors proposed an alternative strategy of privatizing diverse counterfactuals by discretization of continuous features. This technique is closely related to generalization in privacy preserving data mining and is particularly effective against linkage attacks.
@@
K-anonymity using microaggregation has been applied on the training dataset in feature-based XAI by Bozorgpanah & Torra (2024). The study found explanations from non-private and private datasets largely aligned, with minor irregularities observed. The alignment indicated that utility was effectively preserved after privatization. Similarly, Blanco-Justicia et al. (2020) applied microaggregation to build local tree-based surrogate explanations from clusters around an instance to be explained. The method enforced k-anonymity by restricting the cluster size and incorporated shallow trees to enable comprehensibility.
@@
An emerging area of study focusses on providing explanations while protecting privacy in the medical domain, where privacy of patients’ visual data is crucial. The primary aim of such techniques is transformation of private data through removal of identifying features while retaining explanatory evidence. Strategies such as the use of autoencoders for disentanglement of identifiable characteristics (Montenegro & Cardoso, 2024), Siamese network for increasing identity distance between original and privatized images (Montenegro et al., 2021) and latent diffusion models for generating synthetic images Campos et al. (2024) are proposed. The use of lossy compression by pixel sampling is also observed to remove identification information while being post-hoc explainable (Gaudio et al., 2023; Johnson et al., 2022). This approach also has an added advantage of reducing the image size significantly, thus making medical training datasets smaller (Gaudio et al., 2023).
@@
In critical domains, such as healthcare, anonymizing training and query data can assist in protecting identifiable information. However, the applied techniques should preserve the output quality for utility to diverse end-users (Campos et al., 2024). Unlike DP, anonymization techniques lack proven guarantees, however despite DP’s theoretical guarantees, it is unable to scale beyond low resolution image data (Campos et al., 2024). Lossy compression alternatively provides an effective way of privatizing image data, with the benefits of achieving both privacy and explainability while reducing training dataset sizes. It thus enables data sharing with multiple parties in non-private settings (Gaudio et al., 2023) and can serve as an explanation generation method for sensitive image data.
@@
Anonymization techniques, such as k-anonymity, protect privacy of individuals by mitigating re-identification and linkage attacks (Vo et al., 2023). When applying k-anonymity, selecting an appropriate value of k is critical in striking the right balance between accuracy and privacy risk level (Bozorgpanah & Torra, 2024). While higher values of k enhance privacy, explainability may be adversely affected (Berning et al., 2024; Blanco-Justicia et al., 2020). Moreover, the actual level of privacy may also not scale with increasing values of k (Berning et al., 2024). Hence the selection of k that achieves the right trade-off in privacy, explainability and utility is important. Additionally, k-anonymity has limitations such as its dependence on data characteristics, susceptibility to homogeneity attack (Berning et al., 2024), and its vulnerability to privacy leakage when background knowledge is available or diversity is lacking in the private attributes (Goethals et al., 2023). Other techniques such as l-diversity and t-closeness may address some of these challenges, though their applicability to explanations remain unexplored.
@@
The generation of synthetic data for privacy preserving data analysis is explored in previous non-XAI works (Boedihardjo et al., 2023; Liu et al., 2022b). Generating synthetic data that is private, accurate and preserves properties of the true data is a known challenge and NP-hard in the worst case (Ullman & Vadhan, 2011). When models are trained on such data, the explanations through XAI tools are expected to be inherently privacy preserving, hence this approach can be a promising direction in preserving privacy in explainable systems.
@@ 5.4 Perturbation
Perturbation of sensitive data is a widely recognized technique in the field of privacy preserving data publishing (Tran et al., 2024; Yin et al., 2022). When explanations contain sensitive information, obfuscating the contents through perturbations can prevent direct exposure. This technique can also be applied to stem indirect leakage of inferencing sensitive attributes through explanations.
Chen et al. (2022a) proposed a generic privacy preserving mechanism applicable to different XAI types such as feature-based and interpretable surrogates. The proposed method perturbed the decision mapping of an algorithm prior to public release of transparency reports. To mitigate privacy leakage while upholding utility, the authors defined a maximum confidence measure in the inference of sensitive attributes of data subjects and a utility measure in terms of faithfulness. Jeong et al. (2022) applied perturbations on saliency map explanations as a defense mechanism for model inversion in image models. The proposed framework comprised of a two-player minimax game between inversion and noise injector networks, in which the inversion network attempted to reconstruct images from saliency maps and the noise injector perturbed explanations to counter the inversion. The use of multiple evaluation metrics to differentiate between original and reconstructed images facilitated the quantification of the privacy of the defense mechanism.
For the prevention of privacy leakage in XAI, researchers have attempted perturbation of two types of model outputs, namely, predictions and explanations. Adding perturbations to model predictions, such as the strategy of adding noise to output confidence scores used by MemGuard (Jia et al., 2019), is found ineffective in mitigating membership inference through explanations (Liu et al., 2024). Perturbation of explanations is also insufficient in defending against data-free model extraction based on explanations (Yan et al., 2023a). However, the strategy has shown promising results in countering model inversion. The use of perturbation techniques at the explanation interface is also attractive due to its ease of implementation that requires no retraining of the model (Jeong et al., 2022). Nevertheless, large magnitude noise can degrade the usefulness of explanations (Jeong et al., 2022), hence perturbations should be carefully calibrated to minimize any adverse impact on explanation quality.
@@ 5.5 Federated learning
Among the distributed privacy enhancing techniques available in PPML, Federated learning (FL) is an architectural solution (El Mestari et al., 2024) that enables training of local models on user devices and exchange of model parameters with a centralized server that co-ordinates the training of a shared global model (Konečný et al., 2016). It thus enables collaborative learning while keeping users’ private data at the source (Guerra-Manzanares et al., 2023) and mitigates the privacy risk of multiple parties sharing their sensitive data with other parties (El Zein et al., 2024) or a centralized server (Zhu et al., 2022). In horizontal federated learning (HFL), local datasets have the same feature space but contain different samples while vertical federated learning (VFL), involves datasets with different feature spaces but overlaps in samples (Fiosina, 2022).
@@
To address both privacy and explainability in Trustworthy AI, the combination of FL and XAI, i.e., Fed-XAI is suggested (Bárcena et al., 2022; Corcuera Bárcena et al., 2023; López-Blanco et al., 2023) and refers to the federated learning of XAI models. Many approaches of Fed-XAI using HFL and VFL are proposed in literature. Fiosina (2022) used a HFL approach for forecasting taxi trip duration and applied feature-based explainability methods post-hoc. Chen et al. (2022b) used an explainable VFL framework to optimize counterfactual explanations using a representative query distributed on multiple parties. Both setups demonstrate the use of post-hoc explainability tools in a distributed environment, with FL serving as a privacy preserving setup for collaborative learning of sensitive data owned by multiple parties. Fed-XAI architectures have also leveraged interpretable models locally, such as fuzzy rule-based classifiers (Daole et al., 2024), Takagi-Sugeno (Zhu et al., 2022) and Takagi-Sugeno-Kang (Corcuera Bárcena et al., 2023) fuzzy rule-based models. In these setups, interpretability is achieved using underlying explainable-by-design (Corcuera Bárcena et al., 2023) models.
@@
Though FL aids privacy by default, it is prone to reconstruction and inferencing attacks (Mothukuri et al., 2021; Zhang et al., 2024). The sharing of gradients and model parameters, communication mechanism and aggregation process can lead to leakage of privacy of the participating clients (Zhang et al., 2024). Hence researchers have proposed integration of other privacy preserving techniques, such as cryptography, with FL methods. In one such work, Molhoek & Laanen (2024) generated synthetic data on vertically partitioned data in a FL two-party setup. Counterfactuals built from this synthetic data using secure multi-party computation, were ranked and shared with both parties and were found to be resilient to attribute inference. El Zein et al. (2024) proposed a HFL structure using decision tree models, wherein a global decision tree was collaboratively trained by participants and additive secret-sharing was used in aggregation of intermediate results. A VFL technique, Falcon (Wu et al., 2023), utilized a hybrid approach consisting of partially homomorphic encryption (PHE) and additive secret sharing for exchange of intermediate computations. Another setup, Pivot (Wu et al., 2020), proposed as part of Falcon, used threshold partially homomorphic encryption (TPHE) and additive secret sharing to protect privacy of intermediate exchanges. Though these works successfully integrate cryptographic techniques with FL, research has also determined that the use of cryptographic methods in FL reduces the centralized server’s ability to differentiate true model parameters leading to backdoor attacks (Guo et al., 2022). Hence appropriate defense frameworks, such as trust evaluation schemes (Guo et al., 2023), should be incorporated for protection of the FL system from malicious users.
@@
FL enables the training of AI models from diverse, private, and high-quality data (Zhu et al., 2022) located at client systems. It reduces the footprint of user data in the network (Mothukuri et al., 2021) by keeping data at the source and avoids transmission and storage of sensitive information in a centralized location when multiple parties are involved (Wang et al., 2019). Despite its benefits, in its current form FL faces challenges for its risk-free adoption (Mothukuri et al., 2021) including ensuring privacy constraints, merging of local XAI models and dealing with large data streaming that can lead to concept drifts (Bárcena et al., 2022). The introduction of XAI methods in the FL architecture, can also further increase the vulnerability of the system to privacy attacks through explanations. Thus Fed-XAI presently cannot guarantee privacy preservation through XAI components and further research to develop strategies to stem inadvertent privacy leakage through explanations is essential.
@@ 5.6 Other techniques
A client-centric, data-driven approach of generating counterfactuals was proposed by An & Cao (2024) by leveraging previous inferences retrieved by the model user. Due to the generation of counterfactuals locally at the client, the method was shown to be resilient to model extraction while achieving desirable properties such as diversity and succinctness. In another approach to create an interpretable model from a neural network, Marton et al. (2024) described a data-free strategy of distilling the function represented by the model. The method used synthetic data to train a set of neural networks and extracted the parameters to train an Interpretation-Net with an output representation in the form of surrogate decision trees.
Using a knowledge-based approach, Rožanec et al. (2022) applied semantic technologies in the form of domain specific ontology and knowledge graphs, to enhance explanations and describe features on a higher conceptual level. This enabled delinking explanations from features, thus preserving the confidentiality of the underlying model. Further, the integration of feature-based XAI such as LIME, enabled the system to determine features important for predictions. Terziyan & Vitko (2022) also applied semantic techniques to build XAI consisting of decision trees and rules generated from targeted points around the decision boundary of black-box models, without accessing the original training data. Due to the interoperability of semantic rules, the method enabled usage in decentralized setup for collaborative decision making, without individual parties sharing private local data.
These works demonstrate the application of data-free and knowledge-driven techniques in XAI to build privacy-by-design systems for protection of training data and the confidentiality of the model. By disconnecting features from the model and creating abstraction layers for generation of explanations (Rožanec et al., 2022), these strategies are helpful in protecting the underlying assets.
@@
Differential privacy Various DP training algorithms Training data Interpretable (decision trees) Fletcher & Islam (2020)
DP locally linear maps Training data Interpretable (locally linear maps) Harder et al. (2020)
DPSaab Training data Interpretable (FF-CNN) Li et al. (2020a)
DP-EBM Training data Interpretable (EBM) Baek & Chung (2024); Nori et al. (2021)
@@
DP explanation generation Training and query data Feature-based methods using local linear approximations (LIME, etc.) Patel et al. (2022)
@@
Local DP Training data Feature-based methods that aggregate scores (SHAP, etc.) Nguyen et al. (2023)
DP trained SVM Training data Example-based (counterfactuals) Mochaourab et al. (2023)
DP-SGD Training data Feature-based (Grad-CAM) Liu et al. (2024)
@@
DP-RMSProp Training data Example-based (self-influence functions) Cohen & Giryes (2024)
Local DP Training data Feature-based (TreeSHAP) Bozorgpanah & Torra (2024)
Local DP Training data Feature-based (SHAP) Ezzeddine et al. (2024)
@@
DP-WGAN (Wasserstein GAN) Training data Various XAI methods from DALEX framework [^1] DALEX framework is available on https://github.com/modeloriented/dalex Abbasi et al. (2024)
@@
Cryptography Privacy preserving foil trees Training data, model Example-based (contrastive explanations) Veugen et al. (2022)
@@
Additive homomorphic encryption Query data, inference results, model Interpretable (tree-based models) Zhao et al. (2023)
@@
Anonymization Microaggregation (MDAV) Training data, model Interpretable (decision trees) Blanco-Justicia et al. (2020)
@@
Privacy preserving generative model Training data Example-based (case-based) Montenegro et al. (2021)
@@
HeartSpot (lossy compression) Training data Feature-based (saliency maps) Johnson et al. (2022)
@@
Discretization of features (generalization) Training data Example-based (counterfactuals) Vo et al. (2023)
@@
CF-K (k-anonymity of counterfactuals) Training data Example-based (native counterfactuals) Berning et al. (2024); Goethals et al. (2023)
@@
DeepFixCX (lossy compression) Training data Feature-based (saliency maps) Gaudio et al. (2023)
@@
Microaggregation (MDAV) Training data Feature-based (TreeSHAP) Bozorgpanah & Torra (2024)
@@
Disentangled representation learning Training data Example-based (case-based) Montenegro & Cardoso (2024)
@@
Latent diffusion models Training data Example-based (case-based) Campos et al. (2024)
@@
Perturbation GNIME Training and query data Feature-based (saliency maps) Jeong et al. (2022)
@@
Linear-Time Optimal Privacy Scheme Training and query data Various XAI methods (interpretable surrogates, feature-based, etc.) Chen et al. (2022a)
@@
Federated learning Pivot (VFL, additive secret sharing, TPHE) Training data Tree-based models (transparent) Wu et al. (2020)
@@
HFL Training data Feature-based methods (DeepLIFT, integrated gradients, LIME, etc.) Fiosina (2022)
@@
HFL Training data Interpretable (Takagi-Sugeno,Takagi– Sugeno–Kang, fuzzy rule-based classifier) Corcuera Bárcena et al. (2023); Daole et al. (2024); Zhu et al. (2022)
VFL Training data Counterfactuals Chen et al. (2022b)
@@
Falcon (VFL, additive secret sharing, PHE) Explanations and training data Feature-based (LIME) Wu et al. (2023)
@@
PrivaTree (HFL, additive secret sharing) Training data Decision trees (transparent) El Zein et al. (2024)
VFL, SMC, Synthetic data Query data Example-based (counterfactuals) Molhoek & Laanen (2024)
@@
Other techniques Semantic XAI Training data Interpretable (decision trees, semantic rules) Terziyan & Vitko (2022)
@@
Semantic technologies (knowledge graphs, ontologies) Model Feature-based (LIME, etc.) Rožanec et al. (2022)

### v3

@@ 5.3 Anonymization
Anonymization refers to the process of transformation of data (majeedAnonymizationTechniquesPrivacy2021) to obscure the distinctive features of individuals, thus safeguarding their privacy. The process is associated with the removal or modification of direct and quasi-identifiers (majeedAnonymizationTechniquesPrivacy2021), that can uniquely identify individuals. Various methods of anonymization are used in practice, such as k-anonymity, l-diversity, and t-closeness (yinComprehensiveSurveyPrivacypreserving2022). In XAI, different anonymization techniques have been demonstrated in example-based, feature-based, and interpretable methods. Techniques such as disentangled representation learning and lossy compression have been applied on sensitive visual data, such as medical images, to generate privatized explainable-by-design representations.
@@
A dataset is considered k-anonymous if every record is indistinguishable from k-1 other records (sweeneyKanonymityModelProtecting2002), thus providing a measure of the risk of re-identification of records. K-anonymity can be achieved using methods such as suppression and generalization (sweeneyAchievingKanonymityPrivacy2002), which obscure the data to remove identifiable features. Though traditionally this technique is applied to target datasets for protection, goethalsPrivacyIssueCounterfactual2023 proposed its usage on native counterfactuals, that are actual datapoints from the training dataset, for protection against model inversion attack through explanations. This strategy of generating k-anonymous counterfactuals was shown to result in lower information loss and higher validity, outperforming counterfactuals derived from k-anonymized datasets. Further analysis by berningTradeoffPrivacyQuality2024 determined that the effectiveness of k-anonymous counterfactuals is confined to dense areas of the dataset. Its offered privacy protection was also found disproportionate to the value of k. voFeaturebasedLearningDiverse2023 highlighted another limitation, namely, the computational overhead of generating these counterfactuals requiring querying the explainer for a large number of counterfactuals. The authors proposed an alternative strategy of privatizing diverse counterfactuals by discretization of continuous features. This technique is closely related to generalization in privacy preserving data mining and is particularly effective against linkage attacks.
@@
K-anonymity using microaggregation has been applied on the training dataset in feature-based XAI by bozorgpanahExplainableMachineLearning2024. The study found explanations from non-private and private datasets largely aligned, with minor irregularities observed. The alignment indicated that utility was effectively preserved after privatization. Similarly, blanco-justiciaMachineLearningExplainability2020 applied microaggregation to build local tree-based surrogate explanations from clusters around an instance to be explained. The method enforced k-anonymity by restricting the cluster size and incorporated shallow trees to enable comprehensibility.
@@
An emerging area of study focusses on providing explanations while protecting privacy in the medical domain, where privacy of patients’ visual data is crucial. The primary aim of such techniques is transformation of private data through removal of identifying features while retaining explanatory evidence. Strategies such as the use of autoencoders for disentanglement of identifiable characteristics (montenegroAnonymizingMedicalCasebased2024), Siamese network for increasing identity distance between original and privatized images (montenegroPrivacyPreservingGenerativeAdversarial2021) and latent diffusion models for generating synthetic images camposLatentDiffusionModels2024 are proposed. The use of lossy compression by pixel sampling is also observed to remove identification information while being post-hoc explainable (gaudioDeepFixCXExplainablePrivacypreserving2023; johnsonHeartSpotPrivatizedExplainable2022). This approach also has an added advantage of reducing the image size significantly, thus making medical training datasets smaller (gaudioDeepFixCXExplainablePrivacypreserving2023).
@@
In critical domains, such as healthcare, anonymizing training and query data can assist in protecting identifiable information. However, the applied techniques should preserve the output quality for utility to diverse end-users (camposLatentDiffusionModels2024). Unlike DP, anonymization techniques lack proven guarantees, however despite DP’s theoretical guarantees, it is unable to scale beyond low resolution image data (camposLatentDiffusionModels2024). Lossy compression alternatively provides an effective way of privatizing image data, with the benefits of achieving both privacy and explainability while reducing training dataset sizes. It thus enables data sharing with multiple parties in non-private settings (gaudioDeepFixCXExplainablePrivacypreserving2023) and can serve as an explanation generation method for sensitive image data.
@@
Anonymization techniques, such as k-anonymity, protect privacy of individuals by mitigating re-identification and linkage attacks (voFeaturebasedLearningDiverse2023). When applying k-anonymity, selecting an appropriate value of k is critical in striking the right balance between accuracy and privacy risk level (bozorgpanahExplainableMachineLearning2024). While higher values of k enhance privacy, explainability may be adversely affected (berningTradeoffPrivacyQuality2024; blanco-justiciaMachineLearningExplainability2020). Moreover, the actual level of privacy may also not scale with increasing values of k (berningTradeoffPrivacyQuality2024). Hence the selection of k that achieves the right trade-off in privacy, explainability and utility is important. Additionally, k-anonymity has limitations such as its dependence on data characteristics, susceptibility to homogeneity attack (berningTradeoffPrivacyQuality2024), and its vulnerability to privacy leakage when background knowledge is available or diversity is lacking in the private attributes (goethalsPrivacyIssueCounterfactual2023). Other techniques such as l-diversity and t-closeness may address some of these challenges, though their applicability to explanations remain unexplored.
@@
The generation of synthetic data for privacy preserving data analysis is explored in previous non-XAI works (boedihardjoPrivacySyntheticData2023; liuPrivacyPreservingSyntheticData2022). Generating synthetic data that is private, accurate and preserves properties of the true data is a known challenge and NP-hard in the worst case (ullmanPCPsHardnessGenerating2011). When models are trained on such data, the explanations through XAI tools are expected to be inherently privacy preserving, hence this approach can be a promising direction in preserving privacy in explainable systems.
@@ 5.4 Perturbation
Perturbation of sensitive data is a widely recognized technique in the field of privacy preserving data publishing (tranComprehensiveSurveyTaxonomy2024; yinComprehensiveSurveyPrivacypreserving2022). When explanations contain sensitive information, obfuscating the contents through perturbations can prevent direct exposure. This technique can also be applied to stem indirect leakage of inferencing sensitive attributes through explanations.
chenAchievingTransparencyReport2022 proposed a generic privacy preserving mechanism applicable to different XAI types such as feature-based and interpretable surrogates. The proposed method perturbed the decision mapping of an algorithm prior to public release of transparency reports. To mitigate privacy leakage while upholding utility, the authors defined a maximum confidence measure in the inference of sensitive attributes of data subjects and a utility measure in terms of faithfulness. jeongLearningGenerateInversionResistant2022 applied perturbations on saliency map explanations as a defense mechanism for model inversion in image models. The proposed framework comprised of a two-player minimax game between inversion and noise injector networks, in which the inversion network attempted to reconstruct images from saliency maps and the noise injector perturbed explanations to counter the inversion. The use of multiple evaluation metrics to differentiate between original and reconstructed images facilitated the quantification of the privacy of the defense mechanism.
For the prevention of privacy leakage in XAI, researchers have attempted perturbation of two types of model outputs, namely, predictions and explanations. Adding perturbations to model predictions, such as the strategy of adding noise to output confidence scores used by MemGuard (jiaMemGuardDefendingBlackBox2019), is found ineffective in mitigating membership inference through explanations (liuPleaseTellMe2024). Perturbation of explanations is also insufficient in defending against data-free model extraction based on explanations (yanExplanationbasedDatafreeModel2023). However, the strategy has shown promising results in countering model inversion. The use of perturbation techniques at the explanation interface is also attractive due to its ease of implementation that requires no retraining of the model (jeongLearningGenerateInversionResistant2022). Nevertheless, large magnitude noise can degrade the usefulness of explanations (jeongLearningGenerateInversionResistant2022), hence perturbations should be carefully calibrated to minimize any adverse impact on explanation quality.
@@ 5.5 Federated learning
Among the distributed privacy enhancing techniques available in PPML, Federated learning (FL) is an architectural solution (elmestariPreservingDataPrivacy2024) that enables training of local models on user devices and exchange of model parameters with a centralized server that co-ordinates the training of a shared global model (konecnyFederatedLearningStrategies2016). It thus enables collaborative learning while keeping users’ private data at the source (guerra-manzanaresPrivacyPreservingMachineLearning2023) and mitigates the privacy risk of multiple parties sharing their sensitive data with other parties (elzeinPrivaTreeCollaborativePrivacyPreserving2024) or a centralized server (zhuHorizontalFederatedLearning2022). In horizontal federated learning (HFL), local datasets have the same feature space but contain different samples while vertical federated learning (VFL), involves datasets with different feature spaces but overlaps in samples (fiosinaInterpretablePrivacyPreservingCollaborative2022).
@@
To address both privacy and explainability in Trustworthy AI, the combination of FL and XAI, i.e., Fed-XAI is suggested (barcenaFedXAIFederatedLearning2022; corcuerabarcenaEnablingFederatedLearning2023; lopez-blancoFederatedLearningExplainable2023) and refers to the federated learning of XAI models. Many approaches of Fed-XAI using HFL and VFL are proposed in literature. fiosinaInterpretablePrivacyPreservingCollaborative2022 used a HFL approach for forecasting taxi trip duration and applied feature-based explainability methods post-hoc. chenEVFLExplainableVertical2022 used an explainable VFL framework to optimize counterfactual explanations using a representative query distributed on multiple parties. Both setups demonstrate the use of post-hoc explainability tools in a distributed environment, with FL serving as a privacy preserving setup for collaborative learning of sensitive data owned by multiple parties. Fed-XAI architectures have also leveraged interpretable models locally, such as fuzzy rule-based classifiers (daoleTrustworthyAIHeterogeneous2024), Takagi-Sugeno (zhuHorizontalFederatedLearning2022) and Takagi-Sugeno-Kang (corcuerabarcenaEnablingFederatedLearning2023) fuzzy rule-based models. In these setups, interpretability is achieved using underlying explainable-by-design (corcuerabarcenaEnablingFederatedLearning2023) models.
@@
Though FL aids privacy by default, it is prone to reconstruction and inferencing attacks (mothukuriSurveySecurityPrivacy2021; zhangSurveyTrustworthyFederated2024). The sharing of gradients and model parameters, communication mechanism and aggregation process can lead to leakage of privacy of the participating clients (zhangSurveyTrustworthyFederated2024). Hence researchers have proposed integration of other privacy preserving techniques, such as cryptography, with FL methods. In one such work, molhoekSecureCounterfactualExplanations2024 generated synthetic data on vertically partitioned data in a FL two-party setup. Counterfactuals built from this synthetic data using secure multi-party computation, were ranked and shared with both parties and were found to be resilient to attribute inference. elzeinPrivaTreeCollaborativePrivacyPreserving2024 proposed a HFL structure using decision tree models, wherein a global decision tree was collaboratively trained by participants and additive secret-sharing was used in aggregation of intermediate results. A VFL technique, Falcon (wuFalconPrivacyPreservingInterpretable2023), utilized a hybrid approach consisting of partially homomorphic encryption (PHE) and additive secret sharing for exchange of intermediate computations. Another setup, Pivot (wuPrivacyPreservingVertical2020), proposed as part of Falcon, used threshold partially homomorphic encryption (TPHE) and additive secret sharing to protect privacy of intermediate exchanges. Though these works successfully integrate cryptographic techniques with FL, research has also determined that the use of cryptographic methods in FL reduces the centralized server’s ability to differentiate true model parameters leading to backdoor attacks (guoADFLPoisoningAttack2022). Hence appropriate defense frameworks, such as trust evaluation schemes (guoTFLDTTrustEvaluation2023), should be incorporated for protection of the FL system from malicious users.
@@
FL enables the training of AI models from diverse, private, and high-quality data (zhuHorizontalFederatedLearning2022) located at client systems. It reduces the footprint of user data in the network (mothukuriSurveySecurityPrivacy2021) by keeping data at the source and avoids transmission and storage of sensitive information in a centralized location when multiple parties are involved (wangMeasureContributionParticipants2019). Despite its benefits, in its current form FL faces challenges for its risk-free adoption (mothukuriSurveySecurityPrivacy2021) including ensuring privacy constraints, merging of local XAI models and dealing with large data streaming that can lead to concept drifts (barcenaFedXAIFederatedLearning2022). The introduction of XAI methods in the FL architecture, can also further increase the vulnerability of the system to privacy attacks through explanations. Thus Fed-XAI presently cannot guarantee privacy preservation through XAI components and further research to develop strategies to stem inadvertent privacy leakage through explanations is essential.
@@ 5.6 Other techniques
A client-centric, data-driven approach of generating counterfactuals was proposed by anCounterfactualExplanationWill2024 by leveraging previous inferences retrieved by the model user. Due to the generation of counterfactuals locally at the client, the method was shown to be resilient to model extraction while achieving desirable properties such as diversity and succinctness. In another approach to create an interpretable model from a neural network, martonExplainingNeuralNetworks2024 described a data-free strategy of distilling the function represented by the model. The method used synthetic data to train a set of neural networks and extracted the parameters to train an Interpretation-Net with an output representation in the form of surrogate decision trees.
Using a knowledge-based approach, rozanecKnowledgeGraphbasedRich2022 applied semantic technologies in the form of domain specific ontology and knowledge graphs, to enhance explanations and describe features on a higher conceptual level. This enabled delinking explanations from features, thus preserving the confidentiality of the underlying model. Further, the integration of feature-based XAI such as LIME, enabled the system to determine features important for predictions. terziyanExplainableAIIndustry2022 also applied semantic techniques to build XAI consisting of decision trees and rules generated from targeted points around the decision boundary of black-box models, without accessing the original training data. Due to the interoperability of semantic rules, the method enabled usage in decentralized setup for collaborative decision making, without individual parties sharing private local data.
These works demonstrate the application of data-free and knowledge-driven techniques in XAI to build privacy-by-design systems for protection of training data and the confidentiality of the model. By disconnecting features from the model and creating abstraction layers for generation of explanations (rozanecKnowledgeGraphbasedRich2022), these strategies are helpful in protecting the underlying assets.
@@
Differential privacy Various DP training algorithms Training data Interpretable (decision trees) fletcherDecisionTreeClassification2020
DP locally linear maps Training data Interpretable (locally linear maps) harderInterpretableDifferentiallyPrivate2020
DPSaab Training data Interpretable (FF-CNN) liDifferentialPrivacyPreservation2020
DP-EBM Training data Interpretable (EBM) baekDifferentiallyPrivateExplainable2024; noriAccuracyInterpretabilityDifferential2021
@@
DP explanation generation Training and query data Feature-based methods using local linear approximations (LIME, etc.) patelModelExplanationsDifferential2022
@@
Local DP Training data Feature-based methods that aggregate scores (SHAP, etc.) nguyenXRandDifferentiallyPrivate2023
DP trained SVM Training data Example-based (counterfactuals) mochaourabDemonstratorCounterfactualExplanations2023
DP-SGD Training data Feature-based (Grad-CAM) liuPleaseTellMe2024
@@
DP-RMSProp Training data Example-based (self-influence functions) cohenMembershipInferenceAttack2024
Local DP Training data Feature-based (TreeSHAP) bozorgpanahExplainableMachineLearning2024
Local DP Training data Feature-based (SHAP) ezzeddineDifferentialPrivacyAnomaly2024
@@
DP-WGAN (Wasserstein GAN) Training data Various XAI methods from DALEX framework [^1] DALEX framework is available on https://github.com/modeloriented/dalex abbasiFurtherInsightsBalancing2024
@@
Cryptography Privacy preserving foil trees Training data, model Example-based (contrastive explanations) veugenPrivacyPreservingContrastiveExplanations2022
@@
Additive homomorphic encryption Query data, inference results, model Interpretable (tree-based models) zhaoEfficientPrivacypreservingTreebased2023
@@
Anonymization Microaggregation (MDAV) Training data, model Interpretable (decision trees) blanco-justiciaMachineLearningExplainability2020
@@
Privacy preserving generative model Training data Example-based (case-based) montenegroPrivacyPreservingGenerativeAdversarial2021
@@
HeartSpot (lossy compression) Training data Feature-based (saliency maps) johnsonHeartSpotPrivatizedExplainable2022
@@
Discretization of features (generalization) Training data Example-based (counterfactuals) voFeaturebasedLearningDiverse2023
@@
CF-K (k-anonymity of counterfactuals) Training data Example-based (native counterfactuals) berningTradeoffPrivacyQuality2024; goethalsPrivacyIssueCounterfactual2023
@@
DeepFixCX (lossy compression) Training data Feature-based (saliency maps) gaudioDeepFixCXExplainablePrivacypreserving2023
@@
Microaggregation (MDAV) Training data Feature-based (TreeSHAP) bozorgpanahExplainableMachineLearning2024
@@
Disentangled representation learning Training data Example-based (case-based) montenegroAnonymizingMedicalCasebased2024
@@
Latent diffusion models Training data Example-based (case-based) camposLatentDiffusionModels2024
@@
Perturbation GNIME Training and query data Feature-based (saliency maps) jeongLearningGenerateInversionResistant2022
@@
Linear-Time Optimal Privacy Scheme Training and query data Various XAI methods (interpretable surrogates, feature-based, etc.) chenAchievingTransparencyReport2022
@@
Federated learning Pivot (VFL, additive secret sharing, TPHE) Training data Tree-based models (transparent) wuPrivacyPreservingVertical2020
@@
HFL Training data Feature-based methods (DeepLIFT, integrated gradients, LIME, etc.) fiosinaInterpretablePrivacyPreservingCollaborative2022
@@
HFL Training data Interpretable (Takagi-Sugeno,Takagi– Sugeno–Kang, fuzzy rule-based classifier) corcuerabarcenaEnablingFederatedLearning2023; daoleTrustworthyAIHeterogeneous2024; zhuHorizontalFederatedLearning2022
VFL Training data Counterfactuals chenEVFLExplainableVertical2022
@@
Falcon (VFL, additive secret sharing, PHE) Explanations and training data Feature-based (LIME) wuFalconPrivacyPreservingInterpretable2023
@@
PrivaTree (HFL, additive secret sharing) Training data Decision trees (transparent) elzeinPrivaTreeCollaborativePrivacyPreserving2024
VFL, SMC, Synthetic data Query data Example-based (counterfactuals) molhoekSecureCounterfactualExplanations2024
@@
Other techniques Semantic XAI Training data Interpretable (decision trees, semantic rules) terziyanExplainableAIIndustry2022
@@
Semantic technologies (knowledge graphs, ontologies) Model Feature-based (LIME, etc.) rozanecKnowledgeGraphbasedRich2022